OpenStudy (anonymous):

The Coca-Cola company has developed a new soft drink. The art department has been asked to design a label fr this new drink.If the cylindrical can of Coke measures 7 inches tall and has a diameter of 3.5 inches, how much area will the label cover? Round to the nearest tenth.

jimthompson5910 (jim_thompson5910):

the diameter is 3.5 inches, so the radius is d/2 = 3.5/2 = 1.75 inches The lateral surface area is then SA = 2pi*r*h SA = 2pi*1.75*7 SA = 2pi*12.25 SA = 12.25*2pi SA = 24.5pi SA = 24.5(3.14159) SA = 76.968955 SA = 77.0